OpenStudy (anonymous):

What is the quadratic function that is created with roots -10 and -6 and a vertex at (-8, -8)? using the equation f(x) = ax^2 + bx + c

OpenStudy (anonymous):

@SolomonZelman @Nurali @campbell_st

OpenStudy (campbell_st):

ok... so x = -10 is a root then (x + 10) is a factor and x = - 6 is a root (x + 6) is a factor the equation is y = a(x + 10)(x+6) use the point (-8, -8) to find the value of a by substituting x = -8 and y = -8 then you can distribute for the equation

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

so itd be -8 = a(-8 + 10) (-8 + 6) and then a=2 is that right?

OpenStudy (anonymous):

@campbell_st

OpenStudy (campbell_st):

thats correct now substitute it back y = 2(x+10)(x+6) and distribute

OpenStudy (anonymous):

wait, why do i need to do that?

OpenStudy (tkhunny):

Why did you post this question twice? This is also spamming, according to OpenStudy. If you had a question for these volunteers, you should have voiced it here, rather than reposting.